OpenStudy (poetryprincess):

Help Please? A monkey ordered a pizza on January 1st. On January 1st he ate half of the pizza. For the next five days, each day he ate half of what was left of the pizza. How much of the pizza was left after six days?

OpenStudy (anonymous):

Jan 1st eat half of the pizza = 1/2 =.5 2nd eat ( 1- .5 )/2 = .25 3rd eat .25/2 = .125 4th eat .125/2 = .0625 5th eat .0625/2 = .03125 6th eat .03125/2 = .015625 answer is 1 - (add up all together)

OpenStudy (poetryprincess):

Ah, thanks so much! (: